# Post-hoc Study of Climate Microtargeting on Social Media Ads with LLMs: Thematic Insights and Fairness Evaluation

### v3

@@ 1 Introduction
Despite the growing prevalence of microtargeting in climate campaigns, there remains a limited understanding of the specific techniques and linguistic patterns employed to engage different demographic groups. Traditional method Roberts et al. (2019); Vaismoradi et al. (2016, 2013); Braun and Clarke (2012, 2006); Tuckett (2005) often falls short in capturing the nuanced and context-dependent nature of targeted communication. This gap highlights the need for advanced analytical tools that can dissect and interpret the complex language and strategies used in microtargeted climate advertisements. Large language models (LLMs) Brown et al. (2020), with their robust natural language processing (NLP) capabilities, offer a promising solution to this challenge. In this paper, we investigate whether the newly emerged paradigm in NLP; zero-shot prompting of LLMs Brown et al. (2020) and the practice of providing explanations of answers — is better equipped to address those challenges.
@@
Explanations are fundamental to human learning Ahn et al. (1992), as they underscore task principles that facilitate broad generalizations Lombrozo and Carey (2006); Lombrozo (2006). Consider the example text ( “The turbines that provide clean wind energy also create new habitats for fish, keeping the ecosystem healthy." ) from Figure 2(a) for gender identification task. An explanation can elaborate on a brief answer (e.g., male) by connecting it to the broader reasoning process necessary to solve the problem (e.g., “Emphasizing fish habitats may appeal more to men interested in fishing and environmental conservation.”). Thus, explanations enhance understanding by demonstrating how task principles connect questions to their answers.
@@
While LLMs offer powerful capabilities for analyzing and generating text, their widespread use has also highlighted significant challenges related to fairness Kumar et al. (2023); Li et al. (2023b, a); Sag (2023) and bias Lin et al. (2024); Kotek et al. (2023); Fang et al. (2024); Urman and Makhortykh (2023); Esiobu et al. (2023). Research has shown that LLMs, like other machine learning (ML) models, can inherit and even amplify biases present in the data they are trained on Blodgett et al. (2020); Binns (2018). These biases can manifest in various forms, such as differential accuracy across demographic groups Dacon and Liu (2021); Lu et al. (2020); Hitti et al. (2019); Díaz et al. (2018), harmful stereotypes Bianchi et al. (2023); Matthews et al. (2022); Alemany et al. (2022), and discriminatory language patterns Salguero and Espinilla (2018). The implications are particularly concerning in high-stakes applications such as healthcare, finance, and social media, where biased outcomes can perpetuate inequities and undermine trust in artificial intelligence (AI) systems Mehrabi et al. (2021); Buolamwini and Gebru (2018). Addressing these issues requires fairness evaluations using metrics like Demographic Parity, Equal Opportunity, and Predictive Equality to ensure that models perform equitably across all user groups Žliobaitė (2017); Hardt et al. (2016). In the context of this study, examining the fairness of LLMs predictions in demographic targeting is crucial for understanding the microtargeted climate messaging.
@@
In this study, we conduct a post-hoc analysis of climate microtargeting practices on social media by leveraging the power of LLMs (OpenAI’s o1-preview model [^2] https://openai.com/index/learning-to-reason-with-llms/). Post-hoc analysis, typically performed after the main experiment or event, allows us to retrospectively evaluate how effective these campaigns are in targeting specific demographics. Building upon data from previous research by Islam et al. (2023b); Islam and Goldwasser (2024a), we investigate the ability of LLMs to accurately detect targeted messaging based on specific demographic variables, including gender and age group. Additionally, LLMs provide explanations for their classification decisions, offering insights into the thematic and linguistic elements used to engage different audiences. Furthermore, we conduct a comprehensive fairness analysis to identify potential biases in model predictions. Our exploration leads us to the following research questions (RQ), which are crucial for assessing the potential of LLMs to understand microtargeting patterns and provide deeper and more nuanced insights:
@@
  • RQ1: Given a text, can LLMs identify the targeted demographic of the corresponding text accurately and provide the explanation for the reasoning behind the prediction?
@@
  • RQ3: How fair are the LLMs predictions in terms of demographic targeting, and what are the disparities in prediction accuracy and error rates across different demographic groups?
@@ 2 Related Work
Microtargeting on social media platforms such as Meta has been widely studied in the context of political campaigns Islam et al. (2023a); Capozzi et al. (2021); Serrano et al. (2020); Silva et al. (2020); Capozzi et al. (2020) and public health messaging Islam and Goldwasser (2022); Silva and Benevenuto (2021); Mejova and Kalimeri (2020), revealing both the potential benefits and ethical concerns associated with this practice. In the realm of climate communication, microtargeting can be a powerful tool for enhancing message relevance and impact by tailoring content to the specific values and interests of diverse demographic groups Islam et al. (2023b). However, the effectiveness and ethical implications of such targeted messaging remain under-explored, particularly in terms of how different demographic groups are engaged and whether biases are present in the targeting strategies.
Recent advances in LLMs have demonstrated their capability for in-context learning (ICL), significantly enhancing their ability to perform tasks traditionally handled by humans Chowdhery et al. (2023); Kojima et al. (2022); Le Scao et al. (2022); Brown et al. (2020); Gilardi et al. (2023); De Paoli (2023); Dai et al. (2023); Chiang and Lee (2023); Ziems et al. (2024). This progress suggests a strong potential for effectively applying LLMs to our specific task. In the realms of qualitative research (QR) and NLP, innovative methods are being explored to integrate LLMs into Thematic Analysis (TA). Researchers have proposed various frameworks, including an LLMs-in-the-loop approach Islam and Goldwasser (2024b, a); Dai et al. (2023), integrating GPT-3 with expert-designed codebooks Xiao et al. (2023), and developing collaborative interfaces that utilize LLMs for code generation and support in decision-making processes Gao et al. (2023). Other recent work has shown that LLMs can benefit from examples that decompose the reasoning process (can be seen as an explanation), leading to an answer Wei et al. (2022). Despite the impressive capabilities of LLMs, there are concerns about fairness, accountability, and transparency in LLMs predictions which have been highlighted in recent literatures Anthis et al. (2024); Dai et al. (2024); Kumar et al. (2023); Li et al. (2023b); Wu et al. (2023); Bender et al. (2021), emphasizing the need for rigorous evaluation of biases and disparities in model performance across different groups. In this paper, we leverage LLMs to identify targeted demographics and provide explanations of demographic targeting regarding climate-related advertisements on Meta. Besides, we develop a new set of themes and aspects based on those explanations specifically tailored for analyzing messaging. Additionally, we extend previous research by not only focusing on the accuracy of these predictions but also conducting a comprehensive fairness analysis to identify and address potential biases in the model’s performance.
@@ 3 Dataset
For this work, we consider two demographic indicators, i.e., gender and age group. We consider two gender categories, i.e., male vs female. Regarding age group, we consider four age group categories, i.e., young adults whose age range is 18-24, early working age group (25-44), late working age group (45-64), and senior citizens (65+). However, Meta Ad Library API doesn’t provide explicit targeting information beyond ad impressions (views). Note that this paper infers targeting from exclusive impression distributions, which may differ from advertiser intent. In this work, we focus only on ads viewed exclusively by either males or females; ads viewed by both genders were excluded. Therefore, if an ad is viewed solely by males or solely by females, we consider it as a targeted ad. The same logic is applied to age groups. We do not include any ads that overlap across age categories such as young adults, early working-age individuals, late working-age individuals, and senior citizens. By following this methodology, we narrow down the dataset to $227$ ads. Among them, $59$ ads targeting only females and $47$ ads targeting only males. However, we find $25$ ads target only young adults, $82$ ads target only the early working age group, $8$ ads target only the late working age group, and $6$ ads target only senior citizens.
@@ 4 Experimental Setup
To identify the targeted demographics, we use the most recently [^3] September 12, 2024 released by OpenAI, o1-preview model [^4] https://openai.com/index/introducing-openai-o1-preview/. This is a new large language model that uses reinforcement learning (RL) and chain of thought (COT) techniques for complex reasoning, allowing it to think through a detailed internal process before responding to users. We use OpenAI Playground API to run o1-preview by keeping the default parameters. The prompt template for the demographic prediction task with an explanation using LLMs can be found in Fig. 3 in App. A.1. Fig. 2 shows the example prompts for gender prediction. Fig. 4 in App. A.1 shows the example prompts for age group prediction from the climate campaign dataset.
@@ 4.1 Results
Table 1 provides the overall accuracy of the targeted demographic prediction task by LLMs as well as a detailed breakdown of correct and incorrect predictions across each demographic category. LLMs can predict the targeted demographics with an accuracy of $88.55\%$ answering RQ1. LLMs achieve high accuracy in predicting both females ($94.92\%$) and males ($85.10\%$). A small number of females are misclassified as males, and a few males are misclassified as females. Table 1 shows high accuracy for young ($88.00\%$) and early working ($91.46\%$) categories. Performance drops for the late working ($75\%$) age group and significantly for senior ($33.33\%$) categories.
For baseline comparison, we use open sourced LLMs Llama 3 (llama3-70b-8192555https://github.com/meta-llama/llama3) Touvron et al. (2023) and Mistral Large 2 (mistral-large-2407666https://mistral.ai/news/mistral-large-2407/) Jiang et al. (2023). For llama3-70b-8192 and mistral-large-2407, we keep their default parameters. To run llama3-70b-8192, we use Groq API [^7] https://wow.groq.com/. For tokenizer, we use Meta-Llama-3-70B-Instruct from Hugging Face. For running Mistral Large 2, we use Mistral AI API. Besides, we include smaller pre-trained language model BERT Devlin et al. (2019) (bert-base-uncased) for comparison. Furthermore, we compare with simple Logistic Regression (LR) Cox (1958) trained on term frequency–inverse document frequency (tf-idf) features baseline. However, OpenAI’s o1-preview model outperforms the baselines both in gender (macro average F1 score $90.35\%$) and age group (macro average F1 score $71\%$) predictions (Table 2).
@@ 4.2 Statistical Tests
We perform binomial test assessing whether the model’s gender predictions are significantly better than random guessing ($50\%$).
Null hypothesis, $H_{o}$ : probability of correct prediction is $0.5$ (random guessing)
p-value: $\sim 4.37\times 10^{-19}$
Conclusion: Reject $H_{o}$. This means the model’s accuracy is significantly better than chance, with a very strong statistical significance. Table 4 shows the statistical test results from gender subgroup.
@@
Null hypothesis, $H_{o}$: probability of correct prediction is $0.25$ (random guessing)
p-value: $\sim 4.27\times 10^{-45}$
Conclusion: Reject $H_{o}$. The model’s performance is significantly better than random chance. This confirms that the predictions are statistically meaningful. Table 5 shows the statistical test results from age subgroup.
@@ 5.1 Themes and Aspects of Gender Explanations
We prompt LLMs to provide the common theme and aspects under specific theme of the explanations from $40$ correct male predictions and $56$ correct female predictions. The prompt template is shown in Figure 5 App. A.1. We detail the theme of gender explanation and the aspects of the explanation in Table 3. Two experts in NLP and CSS meticulously evaluate the generated themes and aspects derived from the explanations, ensuring their accuracy and alignment with the intended context. The inter-annotator agreement is $0.94$ (almost perfect agreement) using Cohen’s Kappa coefficient Cohen (1960). We detail the theme of the gender explanations and aspects of that explanation in App. A.3.
@@ 5.2 Themes and Aspects of Age Explanations
We prompt LLMs to provide the common theme and aspects under specific theme of the explanations from $22$ correct young adult predictions, $75$ correct early working age group predictions, $6$ correct late working age group predictions, and $2$ correct senior citizen predictions. We detail the theme of the age group explanation and aspects in Table 6. Similarly to gender, the same two NLP and CSS experts review the generated themes and aspects from the explanations to ensure accuracy and contextual alignment. For the age group, the inter-annotator agreement is $0.91$ (almost perfect agreement) using Cohen’s Kappa coefficient. We detail the theme of the age group explanations and aspects of that explanation in App. A.4.
@@ 5.3 Baselines
We add Latent Dirichlet Allocation (LDA) Blei et al. (2003) and BERTopic Grootendorst (2022) as unsupervised techniques for generating topics from LLMs-generated explanations. BERTopic [^8] https://maartengr.github.io/BERTopic/#common is a hierarchical topic modeling and an off-the-shelf neural baseline for clustering texts. BERTopic generates two topics from the prediction explanation of gender and age (Table 13 in App. A.5). We run LDA with four topics and show the top $10$ words in App. A.5 in Table 13. We perform a human evaluation by asking annotators to assess whether the generated themes and topics match with ad’s explanation. Table 12 in App. A.5 shows the human evaluation results on LLM (GPT-4o OpenAI (2024)) assignment, LDA, and BERTopic.
@@ 6 Fairness and Bias Analysis
In this section, we present a comprehensive fairness analysis of the model for gender and age group classifications to answer RQ3. We evaluate the models using established fairness metrics such as Demographic Parity, Equal Opportunity, and Predictive Equality to assess their performance across different groups. By analyzing confusion matrices and classification reports, we identify any disparities in prediction accuracy and error rates between groups. Our analysis aims to identify biases, investigate the underlying reasons for any observed biases. The insights gained from this analysis are critical for guiding future research in developing fair and inclusive algorithms.

### v4

@@ 1 Introduction
Despite the growing prevalence of microtargeting in climate campaigns, there remains a limited understanding of the specific techniques and linguistic patterns employed to engage different demographic groups. Traditional method Roberts et al. (2019); Vaismoradi et al. (2016, 2013); Braun and Clarke (2012, 2006); Tuckett (2005) often falls short in capturing the nuanced and context-dependent nature of targeted communication. This gap highlights the need for advanced analytical tools that can dissect and interpret the complex language and strategies used in microtargeted climate advertisements. Large language models (LLMs) Brown et al. (2020), with their robust natural language processing (NLP) capabilities, offer a promising solution to this challenge. In this paper, we investigate whether the newly emerging paradigm in NLP- zero-shot prompting of LLMs Brown et al. (2020) and the practice of providing explanations of answers are better equipped to address those challenges.
@@
Explanations are fundamental to human learning Ahn et al. (1992), as they underscore task principles that facilitate broad generalizations Lombrozo and Carey (2006); Lombrozo (2006). Consider the example text ( “The turbines that provide clean wind energy also create new habitats for fish, keeping the ecosystem healthy." ) from Figure 2(a) for the gender identification task. An explanation can elaborate on a brief answer (e.g., male) by connecting it to the broader reasoning process necessary to solve the problem (e.g., “Emphasizing fish habitats may appeal more to men interested in fishing and environmental conservation.”). Thus, explanations enhance understanding by demonstrating how task principles connect questions to their answers.
@@
While LLMs offer powerful capabilities for analyzing and generating text, their widespread use has also highlighted significant challenges related to fairness Kumar et al. (2023); Li et al. (2023b, a); Sag (2023) and bias Lin et al. (2024); Kotek et al. (2023); Fang et al. (2024); Urman and Makhortykh (2023); Esiobu et al. (2023). Research has shown that LLMs, like other machine learning (ML) models, can inherit and even amplify biases present in the data on which they are trained Blodgett et al. (2020); Binns (2018). These biases can manifest in various forms, such as differential accuracy across demographic groups Dacon and Liu (2021); Lu et al. (2020); Hitti et al. (2019); Díaz et al. (2018), harmful stereotypes Bianchi et al. (2023); Matthews et al. (2022); Alemany et al. (2022), and discriminatory language patterns Salguero and Espinilla (2018). The implications are particularly concerning in high-stakes applications such as healthcare, finance, and social networks, where biased outcomes can perpetuate inequities and undermine trust in artificial intelligence (AI) systems Mehrabi et al. (2021); Buolamwini and Gebru (2018). Addressing these issues requires fairness evaluations using metrics like Demographic Parity, Equal Opportunity, and Predictive Equality to ensure that models perform equitably across all user groups Žliobaitė (2017); Hardt et al. (2016). In the context of this study, examining the fairness of LLMs’ predictions in demographic targeting is crucial for understanding the microtargeted climate messaging.
@@
In this study, we conduct a post-hoc analysis of climate microtargeting practices on social media by leveraging the power of LLMs (OpenAI’s o1-preview model [^2] https://openai.com/index/learning-to-reason-with-llms/). Post-hoc analysis, typically performed after the main experiment or event, allows us to retrospectively evaluate how effective these campaigns are in targeting specific demographics. Building upon data from previous research by Islam et al. (2023b); Islam and Goldwasser (2024a), we investigate the ability of LLMs to accurately detect targeted messaging based on specific demographic variables, including gender and age group. Additionally, LLMs provide explanations for their classification decisions, offering insights into the thematic and linguistic elements used to engage different audiences. Furthermore, we conduct a comprehensive fairness analysis to identify potential biases in model predictions. Our exploration leads to the following research questions (RQ), which are crucial for assessing the potential of LLMs to understand microtargeting patterns and provide deeper and more nuanced insights:
@@
  • RQ1: Given a text, can LLMs identify the targeted demographic of the corresponding text accurately and provide an explanation for the reasoning behind the prediction?
@@
  • RQ3: How fair are the LLMs’ predictions in terms of demographic targeting, and what are the disparities in prediction accuracy and error rates across different demographic groups?
@@ 2 Related Work
Microtargeting on social media platforms such as Meta has been widely studied in the context of political campaigns Islam et al. (2023a); Capozzi et al. (2021); Serrano et al. (2020); Silva et al. (2020); Capozzi et al. (2020) and public health messaging Islam and Goldwasser (2022); Silva and Benevenuto (2021); Mejova and Kalimeri (2020), revealing both the potential benefits and ethical concerns associated with this practice. In the realm of climate communication, microtargeting can be a powerful tool for enhancing message relevance and impact by tailoring content to the specific values and interests of diverse demographic groups Islam et al. (2023b). However, the effectiveness and ethical implications of such targeted messaging remain underexplored, particularly in terms of how different demographic groups engage and whether biases are present in targeting strategies.
Recent advances in LLMs have demonstrated their capability for in-context learning (ICL), significantly enhancing their ability to perform tasks traditionally handled by humans Chowdhery et al. (2023); Kojima et al. (2022); Le Scao et al. (2022); Brown et al. (2020); Gilardi et al. (2023); De Paoli (2023); Dai et al. (2023); Chiang and Lee (2023); Ziems et al. (2024). This progress suggests a strong potential for effectively applying LLMs to our specific task. In the realms of qualitative research (QR) and NLP, innovative methods are being explored to integrate LLMs into Thematic Analysis (TA) Braun and Clarke (2012). Researchers have proposed various frameworks, including an LLMs-in-the-loop approach Islam and Goldwasser (2024b, a); Dai et al. (2023), integrating GPT-3 with expert-designed codebooks Xiao et al. (2023), and developing collaborative interfaces that utilize LLMs for code generation and support in decision-making processes Gao et al. (2023). Other recent work has shown that LLMs can benefit from examples that decompose the reasoning process (can be seen as an explanation), leading to an answer Wei et al. (2022). Despite the impressive capabilities of LLMs, there are concerns about fairness, accountability, and transparency in LLMs’ predictions, which have been highlighted in recent literature Anthis et al. (2024); Dai et al. (2024); Kumar et al. (2023); Li et al. (2023b); Wu et al. (2023); Bender et al. (2021), emphasizing the need for a rigorous evaluation of biases and disparities in model performance across different groups. In this paper, we leverage LLMs to identify targeted demographics and provide explanations of demographic targeting regarding climate-related advertisements on Meta. Besides, we develop a new set of themes and aspects based on those explanations, specifically tailored for analyzing messaging. Additionally, we extend previous research by not only focusing on the accuracy of these predictions but also conducting a comprehensive fairness analysis to identify and address potential biases in the model’s performance.
@@ 3 Dataset
For this work, we consider two demographic indicators, i.e., gender and age group. We consider two gender categories, i.e., male vs female. Regarding age group, we consider four age group categories, i.e., young adults whose age range is 18-24, early working age group (25-44), late working age group (45-64), and senior citizens (65+). However, the Meta Ad Library API doesn’t provide explicit targeting information beyond ad impressions (views). Note that this paper infers targeting from exclusive impression distributions, which may differ from advertiser intent. In this work, we focus only on ads viewed exclusively by either males or females; ads viewed by both genders were excluded. Therefore, if an ad is viewed solely by males or solely by females, we consider it a targeted ad. The same logic is applied to age groups. We do not include any ads that overlap across age categories, such as young adults, early working-age individuals, late working-age individuals, and senior citizens. By following this methodology, we narrow down the dataset to $227$ ads. Among them, $59$ ads target only females and $47$ ads target only males. However, we find $25$ ads target only young adults, $82$ ads target only the early working age group, $8$ ads target only the late working age group, and $6$ ads target only senior citizens.
@@ 4 Experimental Setup
To identify the targeted demographics, we use the most recently [^3] September 12, 2024 released by OpenAI, o1-preview model [^4] https://openai.com/index/introducing-openai-o1-preview/. This is a new, large language model that uses reinforcement learning (RL) and chain of thought (COT) techniques for complex reasoning, allowing it to think through a detailed internal process before responding to users. We use the OpenAI Playground API to run o1-preview by keeping the default parameters. The prompt template for the demographic prediction task with an explanation using LLMs can be found in Fig. 3 in App. A.1. Fig. 2 shows the example prompts for gender prediction. Fig. 4 in App. A.1 shows the example prompts for age group prediction from the climate campaign dataset.
@@ 4.1 Results
Table 1 provides the overall accuracy of the targeted demographic prediction task by LLMs as well as a detailed breakdown of correct and incorrect predictions across each demographic category. LLMs can predict the targeted demographics with an accuracy of $88.55\%$ answering RQ1. LLMs achieve high accuracy in predicting both females ($94.92\%$) and males ($85.10\%$). A small number of females are misclassified as males, and a few males are misclassified as females. Table 1 shows high accuracy for young ($88.00\%$) and early working ($91.46\%$) categories. Performance drops for the late working ($75\%$) age group and significantly for the senior ($33.33\%$) categories.
For baseline comparison, we use open-sourced LLMs Llama 3 (llama3-70b-8192555https://github.com/meta-llama/llama3) Touvron et al. (2023) and Mistral Large 2 (mistral-large-2407666https://mistral.ai/news/mistral-large-2407/) Jiang et al. (2023). For llama3-70b-8192 and mistral-large-2407, we keep their default parameters. To run llama3-70b-8192, we use Groq API [^7] https://wow.groq.com/. For the tokenizer, we use Meta-Llama-3-70B-Instruct from Hugging Face. For running Mistral Large 2, we use the Mistral AI API. Besides, we include a smaller pre-trained language model BERT Devlin et al. (2019) (bert-base-uncased) for comparison. Furthermore, we compare with simple Logistic Regression (LR) Cox (1958) trained on term frequency–inverse document frequency (tf-idf) features baseline. However, OpenAI’s o1-preview model outperforms the baselines both in gender (macro average F1 score $90.35\%$) and age group (macro average F1 score $71\%$) predictions (Table 2).
@@ 4.2 Statistical Tests
We perform a binomial test assessing whether the model’s gender predictions are significantly better than random guessing ($50\%$).
Null hypothesis, $H_{o}$ : probability of correct prediction is $0.5$ (random guessing), p-value: $\sim 4.37\times 10^{-19}$, Conclusion: Reject $H_{o}$. This means the model’s accuracy is significantly better than chance, with a very strong statistical significance. Table 4 shows the statistical test results from the gender subgroup.
@@
Null hypothesis, $H_{o}$: probability of correct prediction is $0.25$ (random guessing), p-value: $\sim 4.27\times 10^{-45}$, Conclusion: Reject $H_{o}$. The model’s performance is significantly better than random chance. This confirms that the predictions are statistically meaningful. Table 5 shows the statistical test results from the age subgroup.
@@ 5.1 Themes and Aspects of Gender Explanations
We prompt LLMs to provide the common theme and aspects under the specific theme of the explanations from $40$ correct male predictions and $56$ correct female predictions. The prompt template is shown in Figure 5 App. A.1. We detail the theme of gender explanation and the aspects of the explanation in Table 3. Two experts in NLP and CSS meticulously evaluate the generated themes and aspects derived from the explanations, ensuring their accuracy and alignment with the intended context. The inter-annotator agreement is $0.94$ (almost perfect agreement) using Cohen’s Kappa coefficient Cohen (1960). We detail the theme of the gender explanations and aspects of that explanation in App. A.3.
@@ 5.2 Themes and Aspects of Age Explanations
We prompt LLMs to provide the common theme and aspects under the specific theme of the explanations from $22$ correct young adult predictions, $75$ correct early working age group predictions, $6$ correct late working age group predictions, and $2$ correct senior citizen predictions. We detail the theme of the age group explanation and aspects in Table 6. Similarly to gender, the same two NLP and CSS experts review the generated themes and aspects from the explanations to ensure accuracy and contextual alignment. For the age group, the inter-annotator agreement is $0.91$ (almost perfect agreement) using Cohen’s Kappa coefficient. We detail the theme of the age group explanations and aspects of that explanation in App. A.4.
@@ 5.3 Baselines
We add Latent Dirichlet Allocation (LDA) Blei et al. (2003) and BERTopic Grootendorst (2022) as unsupervised techniques for generating topics from LLMs-generated explanations. BERTopic [^8] https://maartengr.github.io/BERTopic/#common is a hierarchical topic modeling and an off-the-shelf neural baseline for clustering texts. BERTopic generates two topics from the prediction explanation of gender and age (Table 13 in App. A.5). We run LDA with four topics and show the top $10$ words in App. A.5 in Table 13. We perform a human evaluation by asking annotators to assess whether the generated themes and topics match the ad’s explanation. Table 12 in App. A.5 shows the human evaluation results on LLM (GPT-4o OpenAI (2024)) assignment, LDA, and BERTopic.
@@ 6 Fairness and Bias Analysis
In this section, we present a comprehensive fairness analysis of the model for gender and age group classifications to answer RQ3. We evaluate the models using established fairness metrics such as Demographic Parity, Equal Opportunity, and Predictive Equality to assess their performance across different groups. By analyzing confusion matrices and classification reports, we identify any disparities in prediction accuracy and error rates between groups. Our analysis aims to identify biases and investigate the underlying reasons for any observed biases. The insights gained from this analysis are critical for guiding future research in developing fair and inclusive algorithms.